OpenStudy (anonymous):

How to find the common ratio of a geometric series? With the first and last term given

OpenStudy (anonymous):

I am trying to solve this: What is the sum of a 6-term geometric series if the first term is 21 and the last term is 352,947 but I don't know how to find its common ratio, help?

OpenStudy (anonymous):

s = a1 / ( r-1) for an infinite GP 352,947 = 21 / ( r-1 ) (r-1 ) = 352947/21 r = 16807+1 = 16808 might be possible i m wrong but this logic help u a lot

OpenStudy (raden):

use the formula : Sn = a1(r^n - 1)/(r - 1)

OpenStudy (anonymous):

Thank you, I got it now :)